OpenStudy (anonymous):

the ratio of men to women in a certain factory is 3 to 4. there are 204 men. how many workers are there ?

OpenStudy (mudkillz):

There should be 272 women. You divide the number of men by 3 then multiply by 4.

OpenStudy (mudkillz):

But you might want to check it! I may be wrong.

OpenStudy (anonymous):

thanks

OpenStudy (mudkillz):

Will you medal?

OpenStudy (mudkillz):

And your welcome.

OpenStudy (anonymous):

sum of ratio=3+4=7 \[\frac{ x }{ 204 }=\frac{ 7 }{ 3 },x=\frac{ 7 }{ 3 }\times 204=?\]